# ColorwAI: Generative Colorways of Textiles through GAN and Diffusion Disentanglement

### v1

@@ 3.2. Disentanglement Methods
In disentanglement, research has empirically shown that given two semantics, there is a hyperplane in the latent space that divides the two factors (Shen et al., 2020). As such, there is a direction $\mathbf{n_{s}}$ in the latent space that manipulates a certain semantic. Given that direction, an image can be manipulated using $\mathbf{z_{1}}\rightarrow\mathbf{z}+\alpha\mathbf{n_{s}}$ to obtain a new image with that direction enhanced. In our case, if we generate an image $\mathbf{x}$ using $\mathbf{z}$ and we wish to make the image bluer, we need to find the direction that disentangles the factor blue $\mathbf{n_{blue}}$ and move positively towards that direction.
To find the traversal directions, the method of InterfaceGAN trains a linear SVM to find the hyperplane that separates each semantic and uses the orthonormal vector exiting the hyperplane as the disentangled direction $\mathbf{n}$ (Shen et al., 2020). Conversely, StyleSpace proposes a method to find the dimensions that deviate the most from the population mean, showing that the variation only present in those dimensions changes the desired semantic (Wu et al., 2021).
3.2.1. ShapleyVec
@@
In ShapleyVec, we build upon these two methods, maintaining their vantage points. First, we use a linear classifier like InterfaceGAN to find the optimal separation boundary between the two classes (images with a given primary color and images without that primary color). Second, we use Shapley values [^2] The method was coined by Lloyd Shapley in 1953 as a method for collaborative game theory, that assigns payouts to players based on their contribution to the total payout. The method is widely used in machine learning interpretability, as it provides a quantitative and comparable value for feature importance. to analyze the contribution of each dimension to the final prediction (Lundberg and Lee, 2017). Inspired by StyleSpace, we use Shapley values to filter out the dimensions that do not contain enough information about the desired feature. The filtering works as follows: we sort the dimensions by their Shapley values in descending order and maintain the first dimensions whose sum of the Shapley values reaches $10<E<$ 1. Shapley values indicate the marginal contribution of each feature, with the total contribution summing to one. Using only the filtered dimensions, we train another linear classifier. We find the final direction $\mathbf{n}$ in a manifold space of $\mathcal{Z}$, where the dimensions that have been filtered out are masked using 00. The orthonormal direction found using the second classifier is used.
4. Experimental Setup
4.0.1. Dataset and Processing
@@
We train our models on the textile collection of the IMET dataset from the Metropolitan Museum of Art, NY, including 20,000 textiles spanning from the sixth century BC to the 21st century across various civilizations. Additionally, we insert 10,000 textile samples from the Victoria and Albert Museum collections. After filtering out black-and-white images, we used 12,918 images for training, 4,324 for validation, and 4,318 for testing. To standardize the images, we remove backgrounds using the language-conditioned Segment Anything Mask (Liu et al., 2023; Kirillov et al., 2023) with the text ’textile’, extracting the largest spanning square within the mask.
4.0.2. Color Annotation
@@
To enable harmonious modifications based on a predominant guiding color, we annotate the main color of each image. Given a latent code $\mathbf{z}$ that generates an image $\mathbf{x}$, we annotate the color $\mathbf{s}$ of $\mathbf{x}$. We rely on CIELab and HSV color encodings, for their alignment with human vision and intuitive representation (Mojsilovic et al., 2000). We extract the color palette from each image, select a suitable color codebook, and quantize the main color in the palette. We empirically find that an adaptation of Bhatia’s method (Bhatia et al., 2004) using CIELab encoding provides the most pleasing and varied palette. We save the palettes and perform adaptive clustering to derive a set of $19$ predominant colors, forming the basis of the disentangled directions. Each image is assigned one of the codebook colors based on HSV similarity weighted $80\%$ on Hue, $10\%$ on Saturation, and $10\%$ on Value, to ensure that the annotations respect the original hue. For more information on the color annotation pipeline refer to Appendix 2.
4.0.3. Disentanglement Evaluation
@@
Inspired by the verification method developed by (Ramesh et al., 2022), we evaluate the disentanglement of each of the color classes using a pseudo-accuracy score. We compare the performance of ShapleyVec to InterfaceGAN and StyleSpace, in StyleGAN and DDPM.
@@
(3) $\alpha_{optimal}=\frac{1}{n}\sum_{i=1}^{N}\alpha*$
@@
(4) $\alpha*=max_{\alpha}[SSIM(\mathbf{z_{0}},\mathbf{z_{\alpha}})\geq 0.75\times SSIM(\mathbf{z_{0}},\mathbf{z_{0}})]$
@@
4.1. Experiments
To find the disentangled direction $\mathbf{n_{c}}$ for each color $c$ in GANs, we annotate $1000$ latent codes $\mathbf{w}\in\mathcal{W}$ which we sample randomly from $\mathcal{N}(0,1)$. We focus on the latent space of StyleGAN2-ADA $\mathcal{W}$. For DDPM, we invert $1000$ random images, which were previously annotated, using deterministic DDIM. We focus on the latent spaces $\mathcal{Z_{t}}$, the UNet input space, and $\mathcal{h_{t}}$, the UNet bottleneck layer, at mixing timestep $t=350$. We experiment with $500100<t<$ 500 and empirically find $350$ to be the best tradeoff between adherence to the original pattern and color variation.
4.1.1. Quantitative Results
@@
DDPM InterfaceGAN 0.253 (0.41) 0.901 (0.15)
DDPM ShapleyVec 0.194 (0.36) 0.883 (0.17)
DDPM StyleSpace 0.247 (0.43) 0.89 (0.16)
Table 1. Results overview.
@@
The results of the evaluation are shown in Table 1. We observe that the standard deviations are generally high due to the great variability in performance among colors, especially for DDPM. Moreover, we see a slight superiority of both ShapleyVec and InterFaceGAN to StyleSpace on StyleGAN. This is to be expected for GANs as StyleSpace is mostly intended to be used on the higher dimensional $\mathcal{W}+$ or $\mathcal{S}$ space of StyleGAN. Furthermore, ShapleyVec marginally surpasses InterFaceGAN in terms of accuracy and relaxed evaluation in StyleGAN, and vice-versa for DDPM. Particularly, ShapleyVec has the best overall p-accuracy, obtaining more precise modifications, InterfaceGAN achieves the best relaxed accuracy, better approximating a general color mood.
@@
Figure 3. Colorways in three colors for three patterns using InterfaceGAN, ShapleyVec, and StyleSpace on StyleGAN and DDPM. The examples are chosen because of their representativity of general observed trends.
[…]
Colorways in three colors for three patterns using InterfaceGAN, ShapleyVec, and StyleSpace on StyleGAN and DDPM. The examples are chosen because of their representativity of general observed trends.
@@
Some visual examples are shown in Figure 3. We observe an astonishing difference between the type of color modification achieved by StyleGAN and by DDPM. The first features more vivid shape and chromatic modifications. The shift to a neighboring pattern, visible in the first and second patterns from the top, appears seamless and visually pleasant, while the attempted transition in the last image of DDPM towards yellow stripes manifests as drastic and incomplete. Furthermore, DDPM maintains more precisely the shape of the original pattern, inserting chromatic modifications in the details. In the first example on the right, we interestingly see that DDPM substitutes the contours with colored flowers in order to change the chromatic profile of the image. This difference may be due to the different setting of the disentanglement: while the modification in StyleGAN takes place prior to any decoding, the modified latent space is inserted in Diffusion only at the mixing step, where the skeleton of the image is already visible.
@@
In terms of visual differences between the methods, the examples show that InterfaceGAN is able to change the color with the smallest shape alterations. In the context of StyleGAN, however, the tonality of the final image is most faithfully achieved by ShapleyVec. For DDPM, we observe a subtle small variation across methods.
4.1.2. Color Representation
@@
An advantage of disentanglement is the creation of an explicit feature representation within the latent space of a model for each semantic. We leverage this property to explore the representation of each color and its features. We also show the confusion matrix of the disentangled directions following the quantitative evaluation. These experiments are displayed in Figure 4. We show the results for ShapleyVec, as it is the only method with meaningful representations in terms of vector similarities and that uses a subset of neurons of the latent space. For clarity, we only show results for a subset of color directions, which cover the spectrum of colors available, sacrificing some shades of brown and blue.
@@
Figure 4. Relationship between color representations for StyleGAN and DDPM using ShapleyVec. The relationships are shown in terms of similarity between the disentangled color vectors, in terms of which dimensions of the latent space encode which color(s), and of how often the modifications using such color directions result in other colors.
[…]
Relationship between color representations for StyleGAN and DDPM using ShapleyVec. The relationships are shown in terms of similarity between the disentangled color vectors, in terms of which dimensions of the latent space encode which color(s), and of how often the modifications using such color directions result in other colors.
@@
The leftmost column pictures the mutual cosine similarities between the feature representations. We observe that the representations within the $W$ space of StyleGAN broadly mirror our perceptual color similarity. In fact, beiges appear close to yellows, in turn, similar to oranges and themselves to browns. Similarly, blues and green are closely related. On the other side, the similarities within DDPM’s $h-space$ are not naturalistic, with the largest similarity appearing between Alpha Blue and Orange-Yellow tones.
@@
The central column figures which neurons in the latent spaces $W$ and $h-space$ are used in the color representations, and the overlap of the neurons across all pairs of colors. Like DDPM’s vectorial representation, the overlaps between colors are not naturalistic. Particularly, the number of overlapping neurons in DDPM mirrors the vector similarities, due to the sparsity of the vectors. In fact, DDPM’s $h-space$ has 18’000 neurons, out of which ShapleyVec uses around 5000 to encode each color. In StyleGAN, we find two colors, Fuji Snow, a white, and Firebird Tail Lights, an orange, that have many common nodes to all other colors, as if they were more general ’color’ representations. In this context, the white could coincide with a black-and-white direction and the orange with a saturation direction.
@@
Lastly, the rightmost column shows the confusion matrices of the modifications using StyleGAN and DDPM. While the confusion in StyleGAN is mostly with neighboring colors in digital spaces, DDPM displays a clear preference towards mutations towards certain colors.
@@
5. Discussion on Computational Creativity
5.0.1. Evaluation
The premise of this framework is to create a stimulating interaction that boosts creativity. It is subjective to assess whether the system provides creative outputs and its influence on the user. While the quantitative evaluation addresses the ’quality’ of how well the re-colorization is achieved, as discussed by Anna Jordanous, SOTA comparisons should accompany but not substitute qualitative evaluation (Jordanous, 2022). The field of Computational Creativity equips us with tools to better frame this question (Ritchie, 2019). Using Ritchie’s terminology, we can distinguish a creative system in a loose sense, thus as carrying out activities considered by society as generally ’creative’, and creativity in a strict sense, an attribute of a specific agent, action, or outcome based on achievement of a certain quality. The colorway creation, pertaining to the design sphere, can be coarsely considered as $creative_{L}$, in the loose sense. What we wish to assess is whether its outputs can be deemed $creative_{S}$, strict sense. To better understand the creativity of the framework, we conducted an expert qualitative evaluation of the framework and its results. This is referred to in the literature as an external evaluation with a symptomatic approach (Ritchie, 2019). That is to say: we consider creativity in the eye of the beholder and inquire into which symptoms of creativity an expert recognizes or not in the system.
The questionnaire is framed within Boden’s creativity evaluation axes: novelty, surprise, acceptability, and quality (Boden, 2004). We assess novelty in P-creativity terms - novelty for the user in question; and surprise in terms of combinatorial, exploratory, and transformational creativity, further explained later (Boden, 2004; Ritchie, 2019). To assess the quality we focus on the aesthetic value of the output and, lastly, we evaluate acceptability with respect to how much the system can be used in practice and, therefore, ’accepted’ by the community.
@@ 5.0.1. Evaluation
Interestingly, the evaluators all displayed a preference for contemporary-looking designs, suggesting a preference for abstract shapes, and saturated and contrasting colors. This highlights a major limitation of the current approach, which is trained on historical textiles and uses nude tones as disentangled directions. The perception of limited novelty and surprise is also influenced by the historical style of the results.
5.0.2. Type of Creativity
Margaret Boden (Boden, 2010) identifies three forms of creativity: combinatorial, exploratory, and transformational creativity. The first is the ability to generate new elements through the combination of existing ones; the second is the generation of new forms through the exploration of conceptual spaces, while the last refers to an abstract transformation.
@@ 5.0.2. Type of Creativity
Within this type of creativity, historicity becomes a value, a framework aimed at being surpassed. The framework allows generating characteristic examples and mutating the colors according to what has been historically seen, whilst allowing for variations. The addition of novel colors, shapes, and other forms of art can further strengthen the ability to learn from the rule and go beyond it. In this sense, following a direction of the latent space posits this type of generation in an advanced stage of Dan Ventura’s Mere Generation steps (Ventura, 2016). In fact, such a process can be considered a filtering step using visual perception, leaving only the inception of a knowledge base as the last missing step towards creativity and away from the highly critiqued ’mere generation’.
6. Conclusion
In conclusion, we have explored the task of ”generative colorway” creation using disentanglement learning. We proposed ShapleyVec, which marginally outperforms other disentanglement methods on StyleGAN and allows the interpretation of the color representation in three ways. We find StyleGAN’s $W$ space to be most aligned with human color perception and to yield the most seamless modifications. We compare the results with disentanglement on DDPM, establishing a shared framework, and identifying a key difference in the editing. While StyleGAN changes ’semantically’ the output, DDPM disentanglement based on the mixing step only allows to ’redirect’ the output toward the new color. Finally, we propose that disentanglement can serve as a method to leverage the exploratory creativity of latent spaces using Margaret Boden’s terminology.
We acknowledge the limitations of this paper which open avenues for future research. Further tests and evaluations should be carried out to be able to assert the superiority ShapleyVec over InterfaceGAN. To enrich the comparison between GAN and Diffusion, more disentanglement methods on DDPM not using the mixing step, such as Asyrp, should be tested. Lastly, to enrich the novelty and surprise of the output, evolutionary art methods can be included in future work.
Appendix A Diffusion Denoising Probabilistic Models
Diffusion models (Ho et al., 2020) generate data by simulating a reverse diffusion process. The core idea is to learn how to reverse a forward diffusion process that gradually adds noise to the data. Starting from pure noise, the reverse process progressively refines the image to the desired outcome.

### v2

@@ 3.2 Disentanglement Methods
In disentanglement, research has empirically shown that given two semantics, there is a hyperplane in the latent space that divides the two factors [41]. As such, there is a direction $\mathbf{n_{s}}$ in the latent space that manipulates a certain semantic. Given that direction, an image can be manipulated using $\mathbf{z_{1}}\rightarrow\mathbf{z}+\alpha\mathbf{n_{s}}$ to obtain a new image with that direction enhanced. In our case, if we generate an image $\mathbf{x}$ using $\mathbf{z}$ and we wish to make the image bluer, we need to find the direction that disentangles the factor blue $\mathbf{n_{blue}}$ and move positively towards that direction.
To find the traversal directions, the method of InterfaceGAN trains a linear SVM to find the hyperplane that separates each semantic and uses the orthonormal vector exiting the hyperplane as the disentangled direction $\mathbf{n}$ [41]. Conversely, StyleSpace proposes a method to find the dimensions that deviate the most from the population mean, showing that the variation only present in those dimensions changes the desired semantic [48].
3.2.1 ShapleyVec
@@
In ShapleyVec, we build upon these two methods, maintaining their vantage points. First, we use a linear classifier like InterfaceGAN to find the optimal separation boundary between the two classes (images with a given primary color and images without that primary color). Second, we use Shapley values [^2] The method was coined by Lloyd Shapley in 1953 as a method for collaborative game theory, that assigns payouts to players based on their contribution to the total payout. The method is widely used in machine learning interpretability, as it provides a quantitative and comparable value for feature importance. to analyze the contribution of each dimension to the final prediction [31]. We filter out the dimensions that do not contain enough information about the desired feature. The filtering works as follows: we sort the dimensions by their Shapley values in descending order and maintain the first dimensions whose sum of the Shapley values reaches $E$ where $10<E<$ 1. Shapley values indicate the marginal contribution of each feature, with the total contribution summing to one. Using only the filtered dimensions, we train another linear classifier. We find the final direction $\mathbf{n}$ in a manifold space of $\mathcal{Z}$, where the dimensions that have been filtered out are masked using 00. The orthonormal direction found using the second classifier is used.
4 Experimental Setup
4.0.1 Dataset and Processing
@@
We train our models on the textile collection of the IMET dataset from the Metropolitan Museum of Art, NY, including 20,000 textiles spanning from the sixth century BC to the 21st century across various civilizations. Additionally, we insert 10,000 textile samples from the Victoria and Albert Museum collections. After filtering out black-and-white images, we used 12,918 images for training, 4,324 for validation, and 4,318 for testing. To standardize the images, we remove backgrounds using the language-conditioned Segment Anything Mask [29, 25] with the text ’textile’, extracting the largest spanning square within the mask.
4.0.2 Color Annotation
@@
To enable modifications based on a predominant guiding color, we annotate the main color of each image. Given a latent code $\mathbf{z}$ that generates an image $\mathbf{x}$, we annotate the color $\mathbf{s}$ of $\mathbf{x}$. We rely on CIELab and HSV color encodings, for their alignment with human vision and intuitive representation [32]. We extract the color palette from each image, select a suitable color codebook, and quantize the main color in the palette. We empirically find that an adaptation of Bhatia’s method [4] using CIELab encoding provides the most pleasing and varied palette. We save the palettes and perform adaptive clustering to derive a set of $19$ predominant colors, forming the basis of the disentangled directions. Each image is assigned one of the codebook colors based on HSV similarity weighted $80\%$ on Hue, $10\%$ on Saturation, and $10\%$ on Value, to ensure that the annotations respect the original hue. For more information on the color annotation pipeline refer to appendices.
4.0.3 Disentanglement Evaluation
@@
Inspired by the verification method developed by [37], we evaluate the disentanglement of each of the color classes using a pseudo-accuracy score. We compare the performance of ShapleyVec to InterfaceGAN and StyleSpace, in StyleGAN and DDM.
@@
$\alpha_{optimal}=\frac{1}{n}\sum_{i=1}^{N}\alpha*$ (3)
@@
$\alpha*=max_{\alpha}[SSIM(\mathbf{z_{0}},\mathbf{z_{\alpha}})\geq 0.75\times SSIM(\mathbf{z_{0}},\mathbf{z_{0}})]$ (4)
@@
4.1 Experiments
To find the disentangled direction $\mathbf{n_{c}}$ for each color $c$ in GANs, we annotate $1000$ latent codes $\mathbf{w}\in\mathcal{W}$ which we sample randomly from $\mathcal{N}(0,1)$. We focus on the latent space of StyleGAN2-ADA $\mathcal{W}$. For DDM, we invert $1000$ random images, which were previously annotated, using deterministic DDIM. We focus on the latent spaces $\mathcal{Z_{t}}$, the UNet input space, and $\mathcal{h_{t}}$, the UNet bottleneck layer, at mixing timestep $t=350$. We experiment with $500100<t<$ 500 and empirically find $350$ to be the best tradeoff between adherence to the original pattern and color variation.
4.1.1 Quantitative Results
@@
DDM InterfaceGAN 0.253 (0.41) 0.901 (0.15)
DDM ShapleyVec 0.194 (0.36) 0.883 (0.17)
DDM StyleSpace 0.247 (0.43) 0.89 (0.16)
Table 1: Results overview.
@@
The results of the evaluation are shown in Table 1. We observe that the standard deviations are generally high due to the great variability in performance among colors, especially for DDM. Moreover, we see a slight superiority of both ShapleyVec and InterFaceGAN to StyleSpace on StyleGAN. This is to be expected for GANs as StyleSpace is mostly intended to be used on the higher dimensional $\mathcal{W}+$ or $\mathcal{S}$ space of StyleGAN. Furthermore, ShapleyVec marginally surpasses InterFaceGAN in terms of accuracy and relaxed evaluation in StyleGAN, and vice-versa for DDM. Particularly, ShapleyVec has the best overall p-accuracy, obtaining more precise modifications, InterfaceGAN achieves the best relaxed accuracy, better approximating a general color mood.
@@
Figure 3: Colorways in three colors for three patterns using InterfaceGAN, ShapleyVec, and StyleSpace on StyleGAN and DDM. The examples are chosen because of their representativity of general observed trends.
@@
Some visual examples are shown in Figure 3. The first observation common to both architectures is that colors are not fully disentangled, and yield minimal shape modifications. While this is a downside in technical terms, the modifications enhance the creative potential of the system. Furthermore, we observe an astonishing difference between the type of color modification achieved by StyleGAN and by DDM. The first features more vivid shape and chromatic modifications. The shift to a neighboring pattern, visible in the first and second patterns from the top, appears seamless and visually pleasant, while the attempted transition in the last image of DDM towards yellow stripes manifests as drastic and incomplete. Furthermore, DDM maintains more precisely the shape of the original pattern, inserting chromatic modifications in the details. This difference may be due to the different setting of the disentanglement: while the modification in StyleGAN takes place prior to any decoding, the modified latent space is inserted in Diffusion only at the mixing step, where the skeleton of the image is already visible.
@@
The examples show that InterfaceGAN is able to change the color with the smallest shape alterations. In the context of StyleGAN, however, the tonality of the final image is most faithfully achieved by ShapleyVec. For DDM, we observe a subtle small variation across methods.
4.1.2 Color Representation
@@
An advantage of disentanglement is the creation of an explicit feature representation within the latent space of a model for each semantic. We leverage this property to explore the representation of each color and its features. We also show the confusion matrix of the disentangled directions following the quantitative evaluation. These experiments are displayed in Figure 4. This highlights a further advantage of ShapleyVec, for which we can inspect representations in terms of vector similarities and the subset of neurons of the latent space.
@@
Figure 4: Relationship between color representations for StyleGAN and DDM using ShapleyVec. The relationships are shown in terms of similarity between the disentangled color vectors, in terms of which dimensions of the latent space encode which color(s), and of how often the modifications using such color directions result in other colors. For clarity, we only show results for a subset of color directions, which cover the spectrum of colors available, sacrificing some shades of brown and blue.
@@
The leftmost column pictures the mutual cosine similarities between the feature representations. We observe that the representations within the $W$ space of StyleGAN broadly mirror our perceptual color similarity. In fact, beiges appear close to yellows, in turn, similar to oranges and themselves to browns. Similarly, blues and green are closely related. On the other side, the similarities within DDM’s $h-space$ are not naturalistic, with the largest similarity appearing between Alpha Blue and orange-yellow tones.
@@
The central column figures which neurons in the latent spaces $W$ and $h-space$ are used in the color representations, and the overlap of the neurons across all pairs of colors. Like DDM’s vectorial representation, the overlaps between colors are not naturalistic. Particularly, the number of overlapping neurons in DDM mirrors the vector similarities, due to the sparsity of the vectors. In fact, DDM’s $h-space$ has 18’000 neurons, out of which ShapleyVec uses around 5000 to encode each color. In StyleGAN, we find two colors, Fuji Snow, a white, and Firebird Tail Lights, an orange, that have many common nodes to all other colors, as if they were more general ’color’ representations. In this context, the white could simulate with a black-and-white direction and the orange with a saturation direction.
@@
Lastly, the rightmost column shows the confusion matrices of the modifications using StyleGAN and DDM. While the confusion in StyleGAN is mostly with neighboring colors in digital spaces, DDM displays a clear preference towards mutations to certain colors.
@@
5 Discussion on Computational Creativity
5.0.1 Evaluation
The premise of this framework is to create a stimulating interaction that boosts creativity. While the quantitative evaluation addresses the quality of the re-colorization, as discussed by Anna Jordanous, SOTA comparisons should accompany but not substitute qualitative evaluation [19]. The field of Computational Creativity equips us with tools to better frame this question [38]. Using Ritchie’s terminology, we can distinguish a creative system in a loose sense, carrying out activities considered by society as generally ’creative’, and creativity in a strict sense, an attribute of a specific agent, action, or outcome based on achievement of a certain quality. The colorway creation, pertaining to the design sphere, can be coarsely considered as $creative_{L}$, in the loose sense. What we wish to assess is whether its outputs can be deemed $creative_{S}$, strict sense. To better understand the creativity of the framework, we conducted an expert qualitative evaluation of the framework and its results. This is referred to in the literature as an external evaluation with a symptomatic approach [38]. We consider creativity in the eye of the beholder and inquire into which symptoms of creativity an expert recognizes or not in the system.
The questionnaire is framed within Boden’s creativity evaluation axes: novelty, surprise, acceptability, and quality [5]. We assess novelty in P-creativity terms - novelty for the user in question; and surprise in terms of combinatorial, exploratory, and transformational creativity, further explained later [5, 38]. To assess the quality we focus on the aesthetic value of the output and, lastly, we evaluate acceptability with respect to how much the system can be used in practice and, therefore, ’accepted’ by the community. Details on the full questionnaire can be found in appendix.
@@ 5.0.1 Evaluation
Interestingly, the evaluators all displayed a preference for contemporary designs, suggesting a preference for abstract shapes, and saturated and contrasting colors. This highlights a major limitation of the current approach, which is trained on historical textiles and uses nude tones as disentangled directions. The perception of limited novelty and surprise is also influenced by the historical style of the results.
5.0.2 Type of Creativity
Margaret Boden [6] identifies three forms of creativity: combinatorial, exploratory, and transformational creativity. The first is the ability to generate new elements through the combination of existing ones; the second is the generation of new forms through the exploration of conceptual spaces, while the last refers to an abstract transformation.
@@ 5.0.2 Type of Creativity
Within this type of creativity, historicity becomes a value, a framework aimed at being surpassed. The framework allows generating characteristic examples and mutating the colors according to what has been historically seen, whilst allowing for variations. The addition of novel colors, shapes, and other forms of art can further strengthen the ability to learn from the rule and go beyond it. In this sense, following a direction of the latent space posits this type of generation in an advanced stage of Dan Ventura’s Mere Generation steps [43]. In fact, such a process can be considered a filtering step using visual perception, leaving only the inception of a knowledge base as the last missing step towards creativity and away from the critiqued ’mere generation’.
6 Conclusion
In conclusion, we have explored the task of "generative colorway" creation using disentanglement learning. We proposed ShapleyVec, which marginally outperforms other disentanglement methods on StyleGAN and allows a more complete interpretation of the color representation in three ways. We find StyleGAN’s $W$ space to be most aligned with human color perception and to yield the most seamless modifications. We compare the results with disentanglement on DDM, establishing a shared framework, and identifying a key visual difference in the edits. While StyleGAN changes ’semantically’ the output, DDM disentanglement based on the mixing step only allows to ’redirect’ the output toward the new color. Finally, we suggest that disentanglement can serve as a method to leverage the exploratory creativity of latent spaces using Margaret Boden’s terminology.
We acknowledge the limitations of this paper which open avenues for future research. Further tests and evaluations should be carried out to be able to assert the superiority ShapleyVec over InterfaceGAN, including a quantitative evaluation for creativity and testing on more samples. To enrich the comparison between GAN and Diffusion, more disentanglement methods on DDM not using the mixing step, such as Asyrp, should be tested. Lastly, to enrich the novelty and surprise of the output, evolutionary art methods, more disentangled factors and data from a larger temporal spectrum can be included in future work.
Acknowledgments
Digital Visual Studies is a project funded by the Max Planck Society. We would like to thank Pepe Ballesteros Zapata and Dario Negueruela del Castillo for the continuous exchange of ideas, Fabrizio Silvestri for the support and for creating contact between the authors; and Valentine Bernasconi and Davide Torre for the invaluable feedback on the manuscript.
Appendix 0.A Diffusion Denoising Models
Diffusion models [17] generate data by simulating a reverse diffusion process. The core idea is to learn how to reverse a forward diffusion process that gradually adds noise to the data. Starting from pure noise, the reverse process progressively refines the image to the desired outcome.